OpenStudy (anonymous):

Okay so i have done this problem about 5 times and keep getting the wrong answer... so can someone explain how to find the slope intercept form co ordinates are (-6,3) and (6,11)

OpenStudy (anonymous):

\[m=y_2-y_1/x_2-x_1\]

OpenStudy (anonymous):

y=mx+b slope intercept form. bigz, do you need help from here?

OpenStudy (anonymous):

Thank you :) simple math mistake i kept making i was simplfying 8/12 to 2/3

OpenStudy (anonymous):

Any time! :) and it happens to us all ;)

OpenStudy (anonymous):

wait... how is 8/12 4/3 and not 2/3. if you divide 8 by 4 and 12 by 4 its 2/3 sorry still slightly confused

OpenStudy (anonymous):

see happens to us all it is 2/3=m

OpenStudy (anonymous):

but then the full equation i get is y=2/3x+7 but online math says that was wrong...

OpenStudy (anonymous):

so now you can use \[y-y_1=m(x-x_1)\] and solve for y

OpenStudy (anonymous):

What is the exact question?

OpenStudy (anonymous):

find the slope intercepet equationof a line that passes through the points (-6,3) and (6,11)

OpenStudy (anonymous):

y-3=2/3(x-(-6) y-3=2/3x +4 y=2/3x+7 ? is still the asnwer i continue to get

OpenStudy (anonymous):

i get: y-11=(2/3)(x-6) y-11=2/3x-4 y=2/3x+7

OpenStudy (anonymous):

okay... :( i must be mis understanding what the online is wanting because i keep getting rregardless of which equation the same asnwer you do.

OpenStudy (anonymous):

can you attach an image of the question or a link?

OpenStudy (anonymous):

The class opens again in a few hours. i will return then and ask the exact phrasing of the question again then. Thanks for the help. I currently believe im mis understanding question and not how to do it :)

OpenStudy (anonymous):

(-6,3) and (6,11) and you have checked the points...

OpenStudy (anonymous):

Yes that was the first thing i did

OpenStudy (anonymous):

ok sorry I couldn't be of more help you have the steps down :)